OpenStudy (savannah_noelle):

which of the following is a solution of x^2+6x=-22? -2+i√10 -4+i√40 -2-i√40 -4-i√10 What are the solutions of 2x^2+3x=-8? -3± i√41/4 -3± i√55/4 -3± i√21/4 -3± i√15/4

OpenStudy (savannah_noelle):

@jim_thompson5910 @Hero Can you help me out please?

OpenStudy (mertsj):

Do you know the quadratic formula?

OpenStudy (mertsj):

Or do you want to solve by completing the square?

OpenStudy (agent0smith):

for \[\Large ax^2+bx+c = 0\] the quadratic formula is\[\Large x=\frac{ -b \pm \sqrt{b^2 - 4ac} }{2a }\]unless you have to complete the square...

OpenStudy (agent0smith):

x^2+6x=-22 first you'll have to get everything on one side, so it's equal to zero

OpenStudy (agent0smith):

^add 22 to both sides

OpenStudy (savannah_noelle):

thank you <333

OpenStudy (anonymous):

u getting it?

OpenStudy (anonymous):

@agent0smith